# Reading Images Like Texts: Sequential Image Understanding in Vision-Language Models

### v1

@@ Abstract
Vision-Language Models (VLMs) have demonstrated remarkable performance across a variety of real-world tasks. However, existing VLMs typically process visual information by serializing images, a method that diverges significantly from the parallel nature of human vision. Moreover, their opaque internal mechanisms hinder both deeper understanding and architectural innovation. Inspired by the dual-stream hypothesis of human vision, which distinguishes the “what” and “where” pathways, we deconstruct the visual processing in VLMs into object recognition and spatial perception for separate study. For object recognition, we convert images into text token maps and find that the model’s perception of image content unfolds as a two-stage process from shallow to deep layers, beginning with attribute recognition and culminating in semantic disambiguation. For spatial perception, we theoretically derive and empirically verify the geometric structure underlying the positional representation in VLMs. Based on these findings, we introduce an instruction-agnostic token compression algorithm based on a plug-and-play visual decoder to improve decoding efficiency, and a RoPE scaling technique to enhance spatial reasoning. Through rigorous experiments, our work validates these analyses, offering a deeper understanding of VLM internals and providing clear principles for designing more capable future architectures. Code is available at https://github.com/Siriuslala/vlm_interp.
@@ 1 Introduction
Generative vision-language models (VLMs) take images and texts as input and generate texts as output. A VLM first obtains the input embeddings of images and texts, which are then concatenated and processed by a language model for next-token prediction. Existing VLMs are typically based on the Transformer architecture (Vaswani et al., 2017) and have achieved strong performance on numerous real-world tasks. However, VLMs are prone to significant hallucination issues, such as incorrectly describing objects or misjudging the spatial relationships between them. These problems not only present challenges for model improvement but also highlight the need for interpretability research, motivating a deeper investigation into the underlying mechanisms behind a model to foster a more comprehensive understanding of it.
The existing research on the interpretability of multimodal models is very limited. Some studies focus on the internal representations of a model, finding the neurons or components (e.g., attention heads) corresponding to the concepts in the real world (Goh et al., 2021; Gandelsman et al., 2023; 2024), or extracting semantic information using logit lens (Neo et al., 2024; Sonia Joseph, 2024; Jiang et al., 2024) or sparse autoencoders (Hugofry, 2024; Daujotas, 2024; Rao et al., 2024). Another line of work employs causal tracing to investigate information storage and transfer (Rajaram et al., 2024; Palit et al., 2023; Basu et al., 2024) or uses blocking-based interventions to study the information flow (Zhang et al., 2024b; 2025) in VLMs. Other approaches provide visualization through heat maps (Selvaraju et al., 2017; Ben Melech Stan et al., 2024; Li et al., 2025). However, current research often overlooks a fundamental distinction between neural networks and human cognition. Specifically, the visual encoder in a VLM is typically a Vision Transformer (ViT) (Dosovitskiy et al., 2020) that adopts the Transformer architecture originally designed for inherently sequential text. When processing an image, a ViT employs raster scan: it partitions the image into patches and flattens them into a one-dimensional sequence. The adjacent patches belonging to the same object may be scattered in different and discontinuous positions in that sequence. In contrast, human visual perception exhibits characteristics of Gestalt cognition (Koffka, 2013), where the brain actively organizes and integrates discontinuous visual signals upon reception, viewing them as a whole. This discrepancy raises critical questions: How do VLMs leverage a 1D image sequence to understand complex 2D concepts and perform tasks like spatial reasoning? And does this cognitive gap between machine and human processing adversely affect VLM performance?
@@ 2.2 Position embedding
In practice, the absolute position embedding could be either pre-calculated by a sinusoidal function (Vaswani et al., 2017) or a set of learnable vectors (Radford et al., 2018; Devlin et al., 2019). Instead of assigning a unique embedding to each position, relative position embedding focuses on the relative distances between different positions. The most commonly used one is RoPE Su et al. (2024), which encodes the absolute position with a rotation matrix and incorporates the explicit relative position dependency in self-attention formulation in each layer. Considering a query $q_{m}=W_{Q}x_{m}=q_{m}^{(0)}+iq_{m}^{(1)}$ and a key $k_{n}=W_{K}x_{n}=k_{n}^{(0)}+ik_{n}^{(1)}(m\neq n)$ in a 2-dimensional input $X\in\mathbb{R}^{n\times 2}$ at two different positions $m$ and $n$, RoPE applies a transformation $f$ to them and the inner product between the query and the key is calculated as follows:
@@
A more detailed formalization and implementation of RoPE is presented in Appendix A.2. In this work, the VLMs we use include LLaVA-1.5 (Liu et al., 2024a), Qwen2 / 2.5-VL (Wang et al., 2024; Bai et al., 2025) and InternVL-2.5 (Chen et al., 2024b). The Qwen2 / 2.5-VL series of models use 2D RoPE in their visual encoders, while all other models use learnable 1D absolute position embedding in ViT. Information about these models can be found in Appendix A.3.
@@ 3.1 Method: logit lens for visual tokens
It is a good way to visualize the geometry of visual representations via representation similarity, while relying solely on the metric in the linear space is insufficient, as shown in Figure 9 and 10. Logit lens is introduced to study the model behaviors from the view of activations (nostalgebraist, 2020). It applies the unembedding matrix $W_{U}\in\mathbb{R}^{D\times|\mathcal{V}|}$ to an activation in order to extract semantic information from it, where $\mathcal{V}$ is the vocabulary of the language model. In VLM, we can apply $W_{U}$ directly to the image representations and get the text tokens corresponding to the image patches:
@@
where $w_{i}^{V}$ is the decoded text token for image patch $p_{i}$, and $H^{l^{\prime}}=(v_{1}^{l^{\prime}},...,v_{N_{V}}^{l^{\prime}})$ is visual part in the output of the $l^{\prime}$-th layer in the LLM. Neo et al. (2024); Jiang et al. (2024) found that the text tokens contain rich semantic information related to their corresponding image areas. To interpret image processing in details, we inspect ViT layer by layer. For the $l$-th ViT layer, we (1) delete the ViT layers after the $l$-th ViT layer and (2) apply logit lens to the outputs of the $l^{\prime}$-th LLM layer. In practice, the value of $l^{\prime}$ is 25 in LLaVA-1.5-7B and 32 in Qwen2.5-VL-7B, as the emergence of the meaningful tokens is the most significant in these layers (see Appendix A.4.2). Thus, we actually create a family of functions $\mathcal{F}=\{f_{l}\}_{l=1}^{L^{V}}$ that map the image patches $P$ to their corresponding text tokens ($f_{l}:P\rightarrow W^{V,l}$) and reflect the dynamic process of object detection via natural language.
@@
To visualize the dynamic process, we propose two concepts: token map and segmentation map. The idea originates from emoji logit lens (Sonia Joseph, 2024) that maps image patches to emojis for visualization. Here a token map is an image where each grid is filled with its corresponding text token. An illustration of the token map is shown in Figure 2. Examples of the token maps in LLaVA-1.5 and Qwen2.5-VL are shown in Figure 15-17 and 18-20, respectively. As for the segmentation map, we create a set of keywords including the names of the main elements in an object (e.g. for bear: {“bear”, “head”, “eye”, “nose”, “paw”}) for each of the main objects in the image. We fill in each grid with the color of object $o_{m}$ if the text token in that grid falls into the keywords set of object $o_{m}$. Details of this process are shown in Algorithm 1. The segmentation maps of LLaVA-1.5 are shown in Figure 3. See Appendix A.4.3 for more examples.
@@ 3.2 Experimental results and analyses
We use the samples from the GQA dataset (Hudson & Manning, 2019). As shown in Figure 3, the geometric structure gradually approaches the shapes of the objects in the original image. We also observe that in the first few token maps, tokens with no practical semantic meaning, such as punctuation or white spaces, account for the majority. From shallow to middle layers, attribute words for common local features (e.g. “fur”, “yellow”) begin to appear. While from middle to deep layers, the attribute words gradually disappear and the representative words (object labels) for specific global object (e.g. “bear”, “rock”) begin to emerge. To quantify this process, we create a set of attribute words $\mathcal{W}^{A,o_{m}}$ and representative words $\mathcal{W}^{R,o_{m}}$ for each object $o_{m}$ in the image, and define the ratio of attribute words as $r_{A}=\frac{1}{N_{V}}\sum_{m=1}^{M}\sum_{w\in\mathcal{W}^{A,o_{m}}}\operatorname{count}(w)$, where $\operatorname{count}(\cdot)$ is the number of occurrences of a token. The ratio of representative words $r_{R}$ is defined likewise.
@@
To further prove this, we evaluate the hallucination of LLaVA-1.5-7B on the POPE (Li et al., 2023) dataset, in which the questions are all in the form of ”Is there a/an [object] …?”. We directly check layer by layer if any of the words in the token maps matches the [object] word. We set the answer to “yes” if any of the text tokens can be found in the synonyms set of the [object] from Lu et al. (2018), and “no” if not. Experimental settings are detailed in Appendix A.4.2. Results show that the accuracy hovers around 50% or less (random guess) until starting to increase at around layer 12, implying that the model has no confidence in what it sees until reaching middle layers. It first perceives discrete low-level features, and then integrate them together and view them as a whole, behaving in a way that follows the Gestalt principles of perceptual (Koffka, 2013) (Appendix A.4.4). By analogy, the model learns to link similar low-level features together via dot-product self-attention (Principle of Similarity and Proximity), and automatically fill in the “gaps” between discontinuous visual tokens via prior knowledge to view them as a whole object (Principle of Closure).
@@ 4.2 Theoretical analyses on 2D RoPE
Therefore, when the model is asked, ”In which direction is A relative to B?”, the components related to object B within object A’s representation becomes the basis for determining the direction. Thus, the model should be able to determine the spatial relationship using only object A’s representation, which will be verified later. However, recent studies on information flow in VLMs (Zhang et al., 2025) indicate that the LLM first attends to specific image regions based on the objects mentioned in the instruction, then aggregates the information into the last token for next-token prediction. Therefore, we further consider the interaction between objects A and B and propose the concept of the direction vector: $v^{r}=h_{o_{S}}^{r}-h_{o_{N}}^{r}(r\in\mathcal{R})$, where $o_{S}$ and $o_{N}$ are the satellite and nucleus, respectively. The subtraction is used for two reasons: (1) Though attention uses weighted sum, subtraction maintains linearity and does not affect our conclusions when comparing different spatial relationships later on. (2) It removes common terms from the two object representations, amplifying their differences and facilitating subsequent experimental analysis. We first compute the direction vectors for “left” and “right” (Eq. 34, 35). By comparison, $v^{left}$ and $v^{right}$ can be written as $c_{1}v_{A}+c_{2}v_{B}-(c_{3}v_{A}+c_{4}v_{B})$ and $c_{1}v_{A}+c_{2}v_{B}+(c_{3}v_{A}+c_{4}v_{B})$, where the terms for distinguishing positions are $c_{3}=(q_{B}^{(0)}k_{A}^{(1)}-q_{B}^{(1)}k_{A}^{(0)})sin(m\theta)$ and $c_{4}=(q_{A}^{(0)}k_{B}^{(1)}-q_{A}^{(1)}k_{B}^{(0)})sin(m\theta)$ (Eq. 36, 37). The opposing vectors in $v^{left}$ and $v^{right}$ indicate the collinearity of their encoded directions. While for $v^{behind}$ the $c_{3}$ and $c_{4}$ terms in it are both zero, it gets the terms $(q_{B}^{(2)}k_{A}^{(3)}-q_{B}^{(3)}k_{A}^{(2)})sin(n\theta)$ and $(q_{A}^{(2)}k_{B}^{(3)}-q_{A}^{(3)}k_{B}^{(2)})sin(n\theta)$, which are absent in $v^{left}$. When comparing $v^{left}$ and $v^{behind}$ (Eq. 38, 39), we find that although they are not strictly orthogonal, their directional information is encoded in different dimensions in the orthogonal subspaces of the $X$ and $Y$ axes. Thus, the model is supposed to separate them easily, which will be verified later. However, the key terms for spatial reasoning (e.g. $c_{3}$, $c_{4}$) are relatively small compared with the common terms, which could potentially limit the model’s ability for spatial reasoning, and a possible solution is presented in Section 5.2.
@@
To verify the above analyses, we use the subset B of the What’s Up dataset (Kamath et al., 2023) (Figure 7(a)) and Qwen2.5-VL-7B. First, we conduct an experiment where we “erase” objects from an image, i.e., we replace the embedding of an object with another and check the change in the model’s prediction of directions. Experimental results are shown in Table 4. We observe that when the embedding of B is replaced with that of A (or vice versa), the token probability of the right answer only drops a little bit (0.909 $\rightarrow$ 0.858 on average). This indicates that an object’s visual embedding contains sufficient information to determine its spatial relationships with other objects. Then, we compute the direction vectors. For each direction in $\mathcal{R}$, we randomly select from the embeddings of A and B and compute $h_{o_{S}}$ and $h_{o_{N}}$ via mean-pooling for 100 times. We visualized the 400 vectors via PCA. Figure 8(a) demonstrates the collinearity and orthogonality of the direction vectors discussed before, while Figure 8(b) shows that the geometry of them is broken without 2D RoPE. Further, we intervene in the visual embeddings of A and B (r=r=“left”) with those in $r^{\prime}=$“right” / “behind”: $V_{o}^{r}]^{\prime}=(1-\alpha)V_{o}^{r}+\alpha\operatorname{Mean}(V_{o}^{r^{\prime}})\cdot 1_{N_{o}}$, where $\alpha$ is the intervention intensity, $N_{o}$ is the number of patches in object $o$, $1_{N_{o}}$ is an all-ones vector in $\mathbb{R}^{N_{o}}$. The direction vector $v^{r}$ will become $(1-\alpha)v^{r}+\alpha\cdot v^{r^{\prime}}$ (see derivations in Equation 51). Results in Figure 8(c) and 8(d) show that the intervention from $V_{o}^{right}$ causes obvious performance drop as $\alpha$ increases, while $V_{o}^{behind}$ does not. This further verifies the geometry of positional information discussed before.

### v2

@@ Abstract
Vision-Language Models (VLMs) have demonstrated remarkable performance across a variety of real-world tasks. However, existing VLMs typically process visual information by serializing images, a method that diverges significantly from the parallel nature of human vision. Moreover, their opaque internal mechanisms hinder both deeper understanding and architectural innovation. Inspired by the dual-stream hypothesis of human vision, which distinguishes the “what” and “where” pathways, we deconstruct the visual processing in VLMs into object recognition and spatial perception for separate study. For object recognition, we convert images into text token maps and find that the model’s perception of image content unfolds as a two-stage process from shallow to deep layers, beginning with attribute recognition and culminating in semantic disambiguation. For spatial perception, we theoretically derive and empirically verify the geometric structure underlying the positional representation in VLMs. Based on these findings, we introduce an instruction-agnostic token compression algorithm based on a plug-and-play visual decoder to improve decoding efficiency, and a RoPE scaling technique to enhance spatial reasoning. Through rigorous experiments, our work validates these analyses, offering a deeper understanding of VLM internals and providing clear principles for designing more capable future architectures  [^1] Code is available at https://github.com/Siriuslala/vlm_interp..
@@ 1 Introduction
Generative vision-language models (VLMs) take images and texts as input and generate texts as output. Existing VLMs are typically based on the Transformer architecture (Vaswani et al., 2017) and have achieved strong performance on numerous real-world tasks. However, VLMs are prone to significant hallucination issues, such as incorrectly describing objects or misjudging the spatial relationships between them. These problems not only present challenges for model improvement but also highlight the need for interpretability research, motivating a deeper investigation into the underlying mechanisms behind a model to foster a more comprehensive understanding of it.
The existing research on the interpretability of multimodal models is very limited. Some studies focus on the internal representations of a model, finding the neurons or components (e.g., attention heads) corresponding to the concepts in the real world (Goh et al., 2021; Gandelsman et al., 2023; 2024), or extracting semantic information using logit lens (Neo et al., 2024; Sonia Joseph, 2024; Jiang et al., 2024) or sparse autoencoders (Hugofry, 2024; Daujotas, 2024; Rao et al., 2024). Another line of work employs causal tracing to investigate information storage and transfer (Rajaram et al., 2024; Palit et al., 2023; Basu et al., 2024) or uses blocking-based interventions to study the information flow (Zhang et al., 2024b; 2025) in VLMs. However, current research often overlooks a fundamental distinction between neural networks and human cognition. Specifically, the visual encoder in a VLM is typically a Vision Transformer (ViT) (Dosovitskiy et al., 2020) that adopts the Transformer architecture originally designed for inherently sequential text. When processing an image, a ViT employs raster scan: it partitions the image into patches and flattens them into a one-dimensional sequence. The adjacent patches belonging to the same object may be scattered in different and discontinuous positions in that sequence. In contrast, human visual perception exhibits characteristics of Gestalt cognition (Koffka, 2013), where the brain actively organizes and integrates discontinuous visual signals upon reception, viewing them as a whole. This discrepancy raises critical questions: How do VLMs leverage a 1D image sequence to understand complex 2D concepts and perform tasks like spatial reasoning? And does this cognitive gap between machine and human processing adversely affect VLM performance?
@@ 2.2 Position embedding
In practice, the absolute position embedding could be either pre-calculated by a sinusoidal function (Vaswani et al., 2017) or a set of learnable vectors (Radford et al., 2018; Devlin et al., 2019). Instead of assigning a unique embedding to each position, relative position embedding focuses on the relative distances between different positions. The most commonly used one is RoPE (Su et al., 2024), which encodes the absolute position with a rotation matrix and incorporates the explicit relative position dependency in self-attention formulation in each layer. Considering a query $q_{m}=W_{Q}x_{m}=q_{m}^{(0)}+iq_{m}^{(1)}$ and a key $k_{n}=W_{K}x_{n}=k_{n}^{(0)}+ik_{n}^{(1)}(m\neq n)$ in a 2-dimensional input $X\in\mathbb{R}^{n\times 2}$ at two different positions $m$ and $n$, RoPE applies a transformation $f$ to them and the inner product between the query and the key is calculated as follows:
@@
A more detailed formalization and implementation of RoPE is presented in Appendix C. In this work, the VLMs we use include LLaVA-1.5 (Liu et al., 2024a), Qwen2 / 2.5-VL (Wang et al., 2024; Bai et al., 2025) and InternVL-2.5 (Chen et al., 2024b). The Qwen2 / 2.5-VL series of models use 2D RoPE in their visual encoders, while all other models use learnable 1D absolute position embedding in ViT. Information about these models can be found in Appendix D.
@@ 3.1 Method: logit lens for visual tokens
It is a good way to visualize the geometry of visual representations via representation similarity, while relying solely on the metric in the linear space is insufficient. This is because the semantics within an individual visual token becomes increasingly complex in deeper ViT layers, due to the interaction between visual tokens via self-attention (see Figure 9 and 10 in Appendix E.1). Thus, the semantics of visual tokens cannot simply be investigated using metrics such as cosine similarity. Logit lens is introduced to study the model behaviors from the view of activations (nostalgebraist, 2020). It applies the unembedding matrix $W_{U}\in\mathbb{R}^{D\times|\mathcal{V}|}$ to an activation in order to extract semantic information from it, where $\mathcal{V}$ is the vocabulary of the language model. In VLM, we can apply $W_{U}$ directly to the image representations and get the text tokens corresponding to the image patches:
@@
where $w_{i}^{V}$ is the decoded text token for image patch $p_{i}$, and $H^{l^{\prime}}=(v_{1}^{l^{\prime}},...,v_{N_{V}}^{l^{\prime}})$ is visual part in the output of the $l^{\prime}$-th layer in the LLM. Neo et al. (2024); Jiang et al. (2024) found that the text tokens contain rich semantic information related to their corresponding image areas. To interpret image processing in details, we inspect ViT layer by layer. For the $l$-th ViT layer, we (1) delete the ViT layers after the $l$-th ViT layer and (2) apply logit lens to the outputs of the $l^{\prime}$-th LLM layer. In practice, the value of $l^{\prime}$ is 25 in LLaVA-1.5-7B and 32 in Qwen2.5-VL-7B, as the emergence of the meaningful tokens is the most significant in these layers (see Appendix E.2). Thus, we actually create a family of functions $\mathcal{F}=\{f_{l}\}_{l=1}^{L^{V}}$ that map the image patches $P$ to their corresponding text tokens ($f_{l}:P\rightarrow W^{V,l}$) and reflect the dynamic process of object detection via natural language.
@@
To visualize the dynamic process, we propose two concepts: token map and segmentation map. The idea originates from emoji logit lens (Sonia Joseph, 2024) that maps image patches to emojis for visualization. Here a token map is an image where each grid is filled with its corresponding text token. An illustration of the token map is shown in Figure 2. Examples of the token maps in LLaVA-1.5 and Qwen2.5-VL are shown in Figure 15-17 and 18-20, respectively. As for the segmentation map, we create a set of keywords including the names of the main elements in an object (e.g. for bear: {“bear”, “head”, “eye”, “nose”, “paw”}) for each of the main objects in the image. We fill in each grid with the color of object $o_{m}$ if the text token in that grid falls into the keywords set of object $o_{m}$. Details of this process are shown in Algorithm 1. The segmentation maps of LLaVA-1.5 are shown in Figure 3. See Appendix E.3 for more examples.
@@ 3.2 Experimental results and analyses
We use the samples from the GQA dataset (Hudson & Manning, 2019). As shown in Figure 3, the geometric structure in segmentation maps gradually approaches the shapes of the objects in the original image. We also observe that in the first few token maps of the earlier layers in ViT, tokens with no practical semantic meaning, such as punctuation or white spaces, account for the majority. From shallow to middle layers, attribute words for common local features (e.g. “fur”, “yellow”) begin to appear. While from middle to deep layers, the attribute words gradually disappear and the representative words (object labels) for specific global object (e.g. “bear”, “rock”) begin to emerge. To quantify this phenomenon, we create a set of attribute words $\mathcal{W}^{A,o_{m}}$ and representative words $\mathcal{W}^{R,o_{m}}$ for each object $o_{m}$ in the image, and define the ratio of attribute words as $r_{A}=\frac{1}{N_{V}}\sum_{m=1}^{M}\sum_{w\in\mathcal{W}^{A,o_{m}}}\operatorname{count}(w)$, where $\operatorname{count}(\cdot)$ is the number of occurrences of a token. The ratio of representative words $r_{R}$ is defined in the same way.
@@
To further prove this, we evaluate the hallucination of LLaVA-1.5-7B on the POPE (Li et al., 2023) dataset, in which the questions are all in the form of “Is there a/an [object] …?”. We directly check layer by layer if any of the words in the token maps matches the [object] word. We set the answer to “yes” if any of the text tokens can be found in the synonyms set of the [object] from Lu et al. (2018), and “no” if not. Experimental settings are detailed in Appendix E.2. Results show that the accuracy hovers around 50% or less (random guess) until starting to increase at around layer 12, implying that the model has no confidence in what it sees until reaching middle layers. It first perceives discrete low-level features, and then integrate them together and view them as a whole, behaving in a way that follows the Gestalt principles of perceptual (Koffka, 2013) (Appendix E.4). By analogy, the model learns to link similar low-level features together via dot-product self-attention (Principle of Similarity and Proximity), and automatically fill in the “gaps” between discontinuous visual tokens via prior knowledge to view them as a whole object (Principle of Closure).
It should be noted that during pre-training, the parameters of the modality connector were adapted to the final layer of the ViT. Therefore, a more rigorous approach would be to train a dedicated modality connector for each ViT layer, though it is laborious to do this for each layer in each model. However, with the existence of the residual stream, every layer in a Transformer contains the output of all previous layers. This means that the last layer of a ViT is essentially a linear superposition of the current layer’s output and the outputs of all previous layers. During VLM pre-training, on a macro level, the output from the last layer of the ViT is aligned with the language space via the modality connector. While from a micro perspective, the modality connector actually processes the last layer of the ViT, along with the output of all preceding layers. Therefore, it should have developed the ability to handle early ViT layers, and thus the reasonableness of our method is guaranteed.
@@ 4.2 Theoretical analyses on 2D RoPE
Therefore, when the model is asked, “In which direction is A relative to B?”, the components related to object B within object A’s representation becomes the basis for determining the direction. Thus, the model should be able to determine the spatial relationship using only object A’s representation, which will be verified later. However, recent studies on information flow in VLMs (Zhang et al., 2025) indicate that the LLM first attends to specific image regions based on the objects mentioned in the instruction, then aggregates the information into the last token for next-token prediction. Therefore, we further consider the interaction between objects A and B and propose the concept of the direction vector: $v^{r}=h_{o_{S}}^{r}-h_{o_{N}}^{r}(r\in\mathcal{R})$, where $o_{S}$ and $o_{N}$ are the satellite and nucleus, respectively. The subtraction is used for two reasons: (1) Though attention uses weighted sum, subtraction maintains linearity and does not affect our conclusions when comparing different spatial relationships later on. (2) It removes common terms from the two object representations, amplifying their differences and facilitating subsequent experimental analysis. We first compute the direction vectors for “left” and “right” (Eq. 34, 35). By comparison, $v^{left}$ and $v^{right}$ can be written as $c_{1}v_{A}+c_{2}v_{B}-(c_{3}v_{A}+c_{4}v_{B})$ and $c_{1}v_{A}+c_{2}v_{B}+(c_{3}v_{A}+c_{4}v_{B})$, where the terms for distinguishing positions are $c_{3}=(q_{B}^{(0)}k_{A}^{(1)}-q_{B}^{(1)}k_{A}^{(0)})sin(m\theta)$ and $c_{4}=(q_{A}^{(0)}k_{B}^{(1)}-q_{A}^{(1)}k_{B}^{(0)})sin(m\theta)$ (Eq. 36, 37). The opposing vectors in $v^{left}$ and $v^{right}$ indicate the collinearity of their encoded directions. While for $v^{behind}$ the $c_{3}$ and $c_{4}$ terms in it are both zero, it gets the terms $(q_{B}^{(2)}k_{A}^{(3)}-q_{B}^{(3)}k_{A}^{(2)})sin(n\theta)$ and $(q_{A}^{(2)}k_{B}^{(3)}-q_{A}^{(3)}k_{B}^{(2)})sin(n\theta)$, which are absent in $v^{left}$. When comparing $v^{left}$ and $v^{behind}$ (Eq. 38, 39), we find that although they are not strictly orthogonal, their directional information is encoded in different dimensions in the orthogonal subspaces of the $X$ and $Y$ axes. Thus, the model is supposed to separate them easily, which will be verified later. However, the key terms for spatial reasoning (e.g. $c_{3}$, $c_{4}$) are relatively small compared with the common terms, which could potentially limit the model’s ability for spatial reasoning, and a possible solution is presented in Section 5.2.
@@
To verify the above analyses, we use the subset B of the What’s Up dataset (Kamath et al., 2023) (Figure 7(a)) and Qwen2.5-VL-7B. First, we conduct an experiment where we “erase” objects from an image, i.e., we replace the embedding of an object with another and check the change in the model’s prediction of directions. Experimental results are shown in Table 4. We observe that when the embedding of B is replaced with that of A (or vice versa), the token probability of the right answer only drops a little bit (0.909 $\rightarrow$ 0.858 on average). This indicates that an object’s visual embedding contains sufficient information to determine its spatial relationships with other objects. Then, we compute the direction vectors. For each direction in $\mathcal{R}$, we randomly select the visual tokens from the embeddings of A and B and compute $h_{o_{S}}$ and $h_{o_{N}}$ via mean-pooling for 100 times. Details can be found in Appendix F.3. We visualized the 400 vectors via PCA. Figure 8(a) demonstrates the collinearity and orthogonality of the direction vectors discussed before, while Figure 8(b) shows that the geometry of them is broken without 2D RoPE. Further, we intervene in the visual embeddings of A and B (where the relationship r=r=“left”) with those in $r^{\prime}=$“right” / “behind” via: $V_{o}^{r}]^{\prime}=(1-\alpha)V_{o}^{r}+\alpha\operatorname{Mean}(V_{o}^{r^{\prime}})\cdot 1_{N_{o}}$, where $\alpha$ is the intervention intensity, $N_{o}$ is the number of patches in object $o$, $1_{N_{o}}$ is an all-ones vector in $\mathbb{R}^{N_{o}}$. After this intervention, the direction vector $v^{r}$ will become $(1-\alpha)v^{r}+\alpha\cdot v^{r^{\prime}}$ (see derivations in Equation 51). Results in Figure 8(c) and 8(d) show that the intervention from $V_{o}^{right}$ causes obvious performance drop as $\alpha$ increases, while $V_{o}^{behind}$ does not. This further verifies the geometry of positional information discussed before.